OpenStudy (anonymous):

Name the property of real numbers illustrated by the equation. -4(x + 3) = -4x - 12 ---I think the correct answer is C--- A. Distributive Property B. Associative Property of Multiplication C. Associative Property of Addition D. Commutative Property of Addition

OpenStudy (texaschic101):

distributive property states : a(b + c) = ab + ac -4(x + 3) = -4x - 12....its A

OpenStudy (texaschic101):

all you doing is taking the number outside of the parenthesis and multiplying by every number in the parenthesis

OpenStudy (anonymous):

Thank You!